### v2

@@ 3 Zero-sum games: punishing strategies
In this section, we provide an overview of relevant results regarding strategies in zero-sum games. They are of interest to construct Nash equilibria with the classical punishment mechanism (previously described in Section 1). Intuitively, this mechanism functions as follows: if some player deviates from the intended outcome of the NE, the other players coordinate as a coalition to prevent the player from having a profitable deviation. The strategy of the coalition used to sabotage the deviating player is called a punishing strategy, and punishing strategies can be obtained via the zero-sum coalition games.
@@ 3.1 Reachability, safety, Büchi and co-Büchi games
Let $\mathcal{G}=(\mathcal{A},\mathsf{Safe}(T))$ be a zero-sum safety game. In $\mathcal{G}$, the winning region $W_{1}(\mathsf{Safe}(T))$ of $\mathcal{P}_{1}$ is a trap for $\mathcal{P}_{2}$: for all $v\in W_{1}(\mathsf{Safe}(T))$, the set $\mathsf{Succ}_{E}(v)\cap W_{1}(\mathsf{Safe}(T))$ is non-empty if $v\in V_{1}$ and $\mathsf{Succ}_{E}(v)\subseteq W_{1}(\mathsf{Safe}(T))$ if $v\in V_{2}$. In particular, strategies of $\mathcal{P}_{1}$ that only select successor vertices that are in $W_{1}(\mathsf{Safe}(T))$ whenever possible are uniformly winning strategies of $\mathcal{P}_{1}$ in $\mathcal{G}$ as they are guaranteed to avoid $T\subseteq W_{2}(\mathsf{Reach}(T))$ from any vertex in $W_{1}(\mathsf{Safe}(T))$. We summarise this property in the following lemma for later.
@@ 3.2 Shortest-path games
We now discuss properties of zero-sum shortest-path games. Let $w\colon E\to\mathbb{N}$ be a weight function and $\mathcal{G}=(\mathcal{A},\mathsf{SPath}^{T}_{w})$ be a zero-sum shortest-path game (fixed for this whole section). First, we note that $\mathcal{G}$ is determined. This can be shown using the determinacy of games with open objectives [21]. An objective $\Omega$ is open if it can be written as a union of sets of continuations of histories, i.e., a union of the form $\bigcup_{h\in\mathcal{H}}\{h\cdot\pi\mid\pi\in\mathsf{Plays}(\mathcal{A})\text{ and }\mathsf{last}(h)=\mathsf{first}(\pi)\}$ for some set $\mathcal{H}\subseteq\mathsf{Hist}(\mathcal{A})$.
@@
The game $\mathcal{G}$ is determined and values in $\mathcal{G}$ are in $\overline{\mathbb{N}}$.
@@
If, for all $\alpha\in\mathbb{N}$, $\mathcal{P}_{2}$ wins from $v_{0}$ in $\mathcal{G}_{\alpha}$, then $\mathsf{val}(v_{0})=+\infty$. Otherwise, $\mathsf{val}(v_{0})$ is the least $\alpha\in\mathbb{N}$ for which $\mathcal{P}_{1}$ wins from $v_{0}$ in $\mathcal{G}_{\alpha}$. This shows that the value of all vertices is defined in $\mathcal{G}$, i.e., $\mathcal{G}$ is determined, and that these values are in $\overline{\mathbb{N}}$.
On the one hand, it can be shown that $\mathcal{P}_{1}$ has a uniformly optimal memoryless strategy. We defer the proof of this statement to A.
@@
In $\mathcal{G}$, $\mathcal{P}_{1}$ has a uniformly optimal memoryless strategy that is uniformly winning in the zero-sum reachability game $(\mathcal{A},\mathsf{Reach}(T))$.
On the other hand, we have illustrated that $\mathcal{P}_{2}$ does not necessarily have an optimal strategy in Example 2.4. We further observe (via the same game) that, in general, there need not be memoryless strategies of $\mathcal{P}_{2}$ that are optimal from all vertices of finite value.
@@
Let $\alpha\in\mathbb{N}$. The construction of $\sigma_{2}^{\alpha}$ can be sketched as follows. On the one hand, $\sigma_{2}^{\alpha}$ is defined so it is winning in the reachability game $(\mathcal{A},\mathsf{Reach}(T))$ from all vertices in $W_{2}(\mathsf{Safe}(T))$. On the other hand, $\sigma_{2}^{\alpha}$ ensures the minimum of $\alpha$ and $\mathsf{val}(v)$ from all vertices $v\in V\setminus W_{2}(\mathsf{Safe}(T))$. Informally, we define $\sigma_{2}^{\alpha}$ as follows. On $W_{2}(\mathsf{Safe}(T))$, we let $\sigma_{2}^{\alpha}$ coincide with a uniformly winning memoryless strategy of $\mathcal{P}_{2}$ in $(\mathcal{A},\mathsf{Reach}(T))$. Outside of $W_{2}(\mathsf{Safe}(T))$, $\sigma_{2}^{\alpha}$ selects a successor vertex such that the sum of the edge weight and the value of the successor vertex is maximal if there is a maximising alternative, and otherwise, $\sigma_{2}^{\alpha}$ selects a successor such that this sum is at least $\alpha$ (which exists because all such sums are in $\overline{\mathbb{N}}$). We show below that defining $\sigma_{2}^{\alpha}$ this way is satisfactory.
@@
We prove that $\sigma_{2}^{\alpha}$ satisfies the claimed properties. First, we remark that any play starting in $W_{2}(\mathsf{Safe}(T))$ consistent with $\sigma_{2}^{\mathsf{Safe}(T)}$ never leaves $W_{2}(\mathsf{Safe}(T))$ – see Lemma 3.2. Property i follows. To establish ii, we show the following property: for any history $h=v_{0}\ldots v_{r}$ that is consistent with $\sigma_{2}^{\alpha}$ such that for all $\ell<r$, $v_{\ell}\notin T$, it holds that $w(h)+\min\{\mathsf{val}(v_{r}),\alpha\}\geq\min\{\mathsf{val}(v_{0}),\alpha\}.$
We proceed by induction on the length of histories. For a history of the form $h=v_{0}$, the property is immediate. We now consider a suitable history $h=v_{0}\ldots v_{r}$ and assume the property holds for $h^{\prime}=v_{0}\ldots v_{r-1}$ by induction (note that $h^{\prime}$ is of the suitable form as well). Let $e=(v_{r-1},v_{r})$.
We discuss two cases depending on whether $\mathsf{val}(v_{r-1})$ is finite and split each case depending on whom controls $v_{r-1}$. We first assume that $\mathsf{val}(v_{r-1})$ is finite. Both players have optimal strategies from any vertex $v$ with finite value. For $\mathcal{P}_{1}$, refer to Theorem 3.5. For $\mathcal{P}_{2}$, by definition of the value, there exists a strategy ensuring $\mathsf{val}(v)+\frac{1}{2}$ from $v$, which is optimal as costs lie in $\overline{\mathbb{N}}$.
@@
Let us now assume that $\mathsf{val}(v_{r-1})=+\infty$. We first consider the case $v_{r-1}\in V_{1}$. Then all successors of $v_{r-1}$ have an infinite value, otherwise $\mathcal{P}_{1}$ could ensure some natural number from $v_{r-1}$ by moving to a successor with finite value and playing optimally from there. We must therefore show that $w(h)+\alpha\geq\min\{\mathsf{val}(v_{0}),\alpha\}$. It follows directly from the induction hypothesis $w(h^{\prime})+\alpha\geq\min\{\mathsf{val}(v_{0}),\alpha\}$ and $w(h)\geq w(h^{\prime})$.
@@ 4.1 Reachability, safety, Büchi and co-Büchi games
We provide characterisations for NE outcomes in games with objectives. The characterisations provided below are analogous to the NE outcome characterisations for finite arenas of [16].
We first consider reachability, Büchi and co-Büchi games: their respective NE outcome characterisations are the same. Let $\mathcal{G}=(\mathcal{A},(\Omega_{i})_{i\in\llbracket{}n\rrbracket{}})$ be a game with reachability, Büchi or co-Büchi objectives. We denote by $W_{i}(\Omega_{i})$ the winning region of the first player of the coalition game $\mathcal{G}_{i}=(\mathcal{A}_{i},\Omega_{i})$, in which $\mathcal{P}_{i}$ is opposed to the other players.
Let $\pi=v_{0}v_{1}v_{2}\ldots\in\mathsf{Plays}(\mathcal{A})$ be a play. Then $\pi$ is an outcome of an NE from $v_{0}$ if and only if, for all $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$, we have $v_{\ell}\notin W_{i}(\Omega_{i})$ for all $\ell\in\mathbb{N}$. On the one hand, if there exist $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$ and $\ell\in\mathbb{N}$ such that $v_{\ell}\in W_{i}(\mathsf{Reach}(T_{i}))$, then $i$ has a profitable deviation by switching to a strategy that is winning from $v_{\ell}$ in $\mathcal{G}_{i}$ when $v_{\ell}$ is reached. Conversely, one constructs a Nash equilibrium as follows. The players follow the play $\pi$, and, if $\mathcal{P}_{i}$ deviates from $\pi$, then all other players conform to a (memoryless) uniformly winning strategy for the second player in the coalition $\mathcal{G}_{i}$ for the objective $\mathsf{Plays}(\mathcal{A})\setminus\Omega_{i}$. We formally state the characterisation below, and provide a correctness proof.
@@
Let $\mathcal{G}=(\mathcal{A},(\Omega_{i})_{i\in\llbracket{}n\rrbracket{}})$ be a game such that, for all $i\in\llbracket{}n\rrbracket{}$, $\Omega_{i}\in\{\mathsf{Reach}(T_{i}),\mathsf{B\ddot{u}chi}(T_{i}),\mathsf{coB\ddot{u}chi}(T_{i})\}$. Let $\pi=v_{0}v_{1}\ldots$ be a play. Then $\pi$ is the outcome of an NE from $v_{0}$ if and only if, for all $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$, $v_{\ell}\notin W_{i}(\Omega_{i})$ for all $\ell\in\mathbb{N}$.
@@
We let $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$. By construction, we obtain that $\pi=\mathsf{Out}(\sigma,v_{0})$. To conclude, we show that $\sigma$ is an NE from $v_{0}$. Players in $\mathsf{SatPl}_{\mathcal{G}}(\pi)$ cannot have profitable deviations. It therefore remains to show that the other players do not have profitable deviations. We let $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$. It suffices to show that for all $\pi^{\prime}\in\mathsf{Plays}(\mathcal{A})$ starting in $v_{0}$, if $\pi^{\prime}$ is consistent with $\sigma_{-i}$, then $\pi^{\prime}\notin\Omega_{i}$. Let $\pi^{\prime}$ be a play starting in $v_{0}$ that is consistent with $\sigma_{-i}$. It follows from the definition of $\sigma$ that $\pi^{\prime}$ can be written as the combination of a prefix $\pi{\ell}$ of $\pi$ and a play $\pi^{\prime\prime}$ consistent with $\tau_{-i}$ starting in the winning region $W_{-i}(\mathsf{Plays}(\mathcal{A})\setminus\Omega_{i})$ of $\mathcal{P}_{-i}$ in $\mathcal{G}_{i}$. In particular, we have $\pi^{\prime\prime}\notin\Omega_{i}$. If $\Omega_{i}$ is a reachability objective, the target of $\mathcal{P}_{i}$ does not occurs in $\pi{\ell}$ (as $\pi\notin\Omega_{i}$) and does not occur in $\pi^{\prime\prime}$ either (as $\pi^{\prime\prime}\notin\Omega_{i}$). If $\Omega_{i}$ is a Büchi or co-Büchi objective, we have $\pi^{\prime}\in\Omega_{i}$ if and only if $\pi^{\prime\prime}\in\Omega_{i}$. In both cases, we obtain that $\pi^{\prime}\notin\Omega_{i}$. This shows that $\mathcal{P}_{i}$ does not have a profitable deviation, and thus that $\sigma$ is an NE from $v_{0}$.
For safety objectives, we have to slightly adapt the characterisation of Theorem 4.1. If a safety objective is violated by a play, this violation is witnessed by a prefix of the play. It follows that there can be no profitable deviations for some $\mathcal{P}_{i}$ that occurs at a point of an outcome after a vertex unsafe for $\mathcal{P}_{i}$ is traversed. In other words, in a safety game $\mathcal{G}=(\mathcal{A},(\mathsf{Safe}(T_{i}))_{i})$, a play $\pi$ is an NE outcome if and only if, for all $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$, all vertices of $\pi$ that occur before $T_{i}$ is visited are outside of $W_{i}(\mathsf{Safe}(T_{i}))$. As the proof of this claim is analogous to that of Theorem 4.1, we omit it.

### v3

@@ 3 Zero-sum games: punishing strategies
In this section, we provide an overview of relevant results regarding strategies in zero-sum games. They are of interest to construct Nash equilibria with the classical punishment mechanism (previously described in Section 1). Intuitively, this mechanism functions as follows: if some player deviates from the intended outcome of the NE, the other players coordinate as a coalition to ensure that no deviation is profitable by punishing a player who deviates. The strategy of the coalition used to sabotage the deviating player is called a punishing strategy, and punishing strategies can be obtained via the zero-sum coalition games.
@@ 3.1 Reachability, safety, Büchi and co-Büchi games
Let $\mathcal{G}=(\mathcal{A},\mathsf{Safe}(T))$ be a zero-sum safety game. In $\mathcal{G}$, the winning region $W_{1}(\mathsf{Safe}(T))$ of $\mathcal{P}_{1}$ is a trap for $\mathcal{P}_{2}$: for all $v\in W_{1}(\mathsf{Safe}(T))$, the set $\mathsf{Succ}_{E}(v)\cap W_{1}(\mathsf{Safe}(T))$ is non-empty if $v\in V_{1}$ and $\mathsf{Succ}_{E}(v)\subseteq W_{1}(\mathsf{Safe}(T))$ if $v\in V_{2}$. In particular, strategies of $\mathcal{P}_{1}$ that only select successor vertices that are in $W_{1}(\mathsf{Safe}(T))$ whenever possible are uniformly winning strategies of $\mathcal{P}_{1}$ in $\mathcal{G}$ as they are guaranteed to avoid $T\subseteq W_{2}(\mathsf{Reach}(T))$ from any vertex in $W_{1}(\mathsf{Safe}(T))$. We summarise this classical property in the following lemma for later.
@@ 3.2 Shortest-path games
We now discuss properties of zero-sum shortest-path games. Let $w\colon E\to\mathbb{N}$ be a weight function and $\mathcal{G}=(\mathcal{A},\mathsf{SPath}^{T}_{w})$ be a zero-sum shortest-path game (fixed for this whole section). First, we note that $\mathcal{G}$ is determined. This can be shown using the determinacy of games with open objectives [21]. An objective $\Omega$ is open if it can be written as a union of sets of continuations of histories, i.e., a union of the form $\bigcup_{h\in\mathcal{H}}\{h\cdot\pi\mid\pi\in\mathsf{Plays}(\mathcal{A})\text{ and }\mathsf{last}(h)=\mathsf{first}(\pi)\}$ for some set $\mathcal{H}\subseteq\mathsf{Hist}(\mathcal{A})$. The following argument also shows that $\mathcal{P}_{1}$ has optimal strategies from all vertices and $\mathcal{P}_{2}$ has optimal strategies from vertices with a finite value.
@@
The shortest-path game $\mathcal{G}$ is determined. For all $v_{0}\in V$, $\mathcal{P}_{1}$ has an optimal strategy from $v_{0}$ and, if $\mathsf{val}(v_{0})<+\infty$, then $\mathcal{P}_{2}$ has an optimal strategy from $v_{0}$.
@@
Let $v_{0}\in V$. We consider two cases. First, assume that for all $\alpha\in\mathbb{N}$, $\mathcal{P}_{2}$ wins from $v_{0}$ in $\mathcal{G}_{\alpha}$. It follows that $\mathsf{val}(v_{0})=+\infty$ and that all strategies of $\mathcal{P}_{1}$ are optimal from $v_{0}$. We now assume that $\mathcal{P}_{1}$ wins from $v_{0}$ in $\mathcal{G}_{\alpha}$ for some $\alpha\in\mathbb{N}$. We have $\mathsf{val}(v_{0})=\min\{\alpha\in\mathbb{N}\mid\mathcal{P}_{1}\text{ wins from }v_{0}\text{ in }\mathcal{G}_{\alpha}\}$, as both players have a strategy ensuring this number from $v_{0}$ by the above (and the fact that all strategies of $\mathcal{P}_{2}$ ensure 0 from any vertex in $\mathcal{G}$). In particular, both players have optimal strategies in $\mathcal{G}$ from $v_{0}$.
On the one hand, it can be shown that $\mathcal{P}_{1}$ has a uniformly optimal memoryless strategy. We defer the proof of this statement to Appendix A.
@@
In the shortest-path $\mathcal{G}$, $\mathcal{P}_{1}$ has a uniformly optimal memoryless strategy that is uniformly winning in the zero-sum reachability game $(\mathcal{A},\mathsf{Reach}(T))$.
On the other hand, we have illustrated that $\mathcal{P}_{2}$ does not necessarily have an optimal strategy from states with an infinite value in Example 2.4. We further observe (via the same game) that, in general, there need not be memoryless strategies of $\mathcal{P}_{2}$ that are optimal from all vertices of finite value.
@@
Let $\alpha\in\mathbb{N}$. We define $\sigma_{2}^{\alpha}$ as follows. On $W_{2}(\mathsf{Safe}(T))$, we let $\sigma_{2}^{\alpha}$ coincide with a uniformly winning memoryless strategy of $\mathcal{P}_{2}$ in $(\mathcal{A},\mathsf{Reach}(T))$. For all $v\in V_{2}\setminus W_{2}(\mathsf{Safe}(T))$, we $\sigma_{2}^{\alpha}(v)$ be a successor $v^{\prime}$ of $v$ maximising $w(v,v^{\prime})+\mathsf{val}(v^{\prime})$ if there exists one, and, otherwise let $\sigma_{2}^{\alpha}(v)$ be a successor $v^{\prime}$ of $v$ such that $w(v,v^{\prime})+\mathsf{val}(v^{\prime})\geq\alpha$. This ensures that from any vertex with an infinite value outside of $W_{2}(\mathsf{Safe}(T))$, $\mathcal{P}_{2}$ ensures a cost of at least $\alpha$. In particular, from any vertex $v$ with $\mathsf{val}(v)<\alpha$, $\mathcal{P}_{1}$ cannot exploit the decisions of $\mathcal{P}_{2}$ in infinite-value vertices to obtain a cost less than $\mathsf{val}(v)$ from $v$ (as illustrated in Example 2.4). We formally establish the required properties of $\sigma_{2}^{\alpha}$ in the following proof.
@@
We prove that $\sigma_{2}^{\alpha}$ satisfies the claimed properties. To show i, it suffices to prove that any play starting in $W_{2}(\mathsf{Safe}(T))$ consistent with $\sigma_{2}^{\mathsf{Safe}(T)}$ never leaves $W_{2}(\mathsf{Safe}(T))$. We can prove this by contradiction. Assume that there there exists $v\in W_{2}(\mathsf{Safe}(T))$ such that $\sigma_{2}^{\mathsf{Safe}(T)}(v)\notin W_{2}(\mathsf{Safe}(T))$. By determinacy of safety games, $\mathcal{P}_{1}$ can force a visit to $T$ from $\sigma_{2}^{\mathsf{Safe}(T)}(v)$. This implies that $\sigma_{2}^{\mathsf{Safe}(T)}$ is not winning from $v$, contradicting the fact that $\sigma_{2}^{\mathsf{Safe}(T)}$ is uniformly winning.
To establish ii, we show the following property: for any history $h=v_{0}\ldots v_{r}$ that is consistent with $\sigma_{2}^{\alpha}$ such that for all $\ell<r$, $v_{\ell}\notin T$, it holds that $w(h)+\min\{\mathsf{val}(v_{r}),\alpha\}\geq\min\{\mathsf{val}(v_{0}),\alpha\}.$
We proceed by induction on the length of histories. For all $v_{0}\in V$, the property is immediate for the history $h=v_{0}$. We now consider a suitable history $h=v_{0}\ldots v_{r}$ with $r\geq 1$ and assume the property holds for $h^{\prime}=v_{0}\ldots v_{r-1}$ by induction (note that $h^{\prime}$ is of the suitable form as well). Let $e=(v_{r-1},v_{r})$.
We discuss two cases depending on whether $\mathsf{val}(v_{r-1})$ is finite and split each case depending on whom controls $v_{r-1}$. We first assume that $\mathsf{val}(v_{r-1})$ is finite. By Lemma 3.3, both players have optimal strategies from any vertex with a finite value.
@@
Let us now assume that $\mathsf{val}(v_{r-1})=+\infty$. We first consider the case $v_{r-1}\in V_{1}$. Then all successors of $v_{r-1}$ have an infinite value, otherwise $\mathcal{P}_{1}$ could ensure some natural number from $v_{r-1}$ by moving to a successor with finite value and playing optimally from there. We must therefore show that $w(h)+\alpha\geq\min\{\mathsf{val}(v_{0}),\alpha\}$, which follows directly from $\alpha\geq\min\{\mathsf{val}(v_{0}),\alpha\}$.
@@ 4.1 Reachability, safety, Büchi and co-Büchi games
We provide characterisations for NE outcomes in games with objectives. The characterisations provided in this section are analogous to the NE outcome characterisations for finite arenas of [16]. These characterisations can be proven in the same way in an arbitrary arena as they can be in finite arenas; we include a proof below to illustrate the construction of NEs via a punishment mechanism.
First, we consider games with reachability, Büchi and co-Büchi objectives. Let $\mathcal{G}=(\mathcal{A},(\Omega_{i})_{i\in\llbracket{}n\rrbracket{}})$ be a game with $\Omega_{i}\in\{\mathsf{Reach}(T_{i}),\mathsf{B\ddot{u}chi}(T_{i}),\mathsf{coB\ddot{u}chi}(T_{i})\}$ for all $i\in\llbracket{}n\rrbracket{}$. We denote by $W_{i}(\Omega_{i})$ the winning region of the first player of the coalition game $\mathcal{G}_{i}=(\mathcal{A}_{i},\Omega_{i})$, in which $\mathcal{P}_{i}$ is opposed to the other players.
Let $\pi=v_{0}v_{1}v_{2}\ldots\in\mathsf{Plays}(\mathcal{A})$ be a play. The play $\pi$ is an NE outcome if and only if the players whose objectives are not satisfied cannot enforce their objective from any vertex along $\pi$, i.e., if for all $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$, $v_{\ell}\notin W_{i}(\Omega_{i})$ for all $\ell\in\mathbb{N}$. On the one hand, if some player whose objective is not satisfied can enforce it from some vertex $v_{\ell}$ along $\pi$, then they have a profitable deviation by switching to a winning strategy of their coalition game from $v_{\ell}$; the resulting outcome is winning for this player as we consider reachability, Büchi and co-Büchi objectives. Conversely, one constructs a Nash equilibrium as follows. The players follow the play $\pi$, and, if $\mathcal{P}_{i}$ deviates from $\pi$, then all other players conform to a (memoryless) uniformly winning strategy for the second player in the coalition $\mathcal{G}_{i}$ for the objective $\mathsf{Plays}(\mathcal{A})\setminus\Omega_{i}$. This ensures no player has a profitable deviation.
We formally state the characterisation below. The following statement differs slightly from its finite-arena counterpart in [16]: we consider games in which players need not have objectives of the same type, whereas [16] considers games in which either all players have reachability objectives or all players have prefix-independent objectives (which includes Büchi and co-Büchi objectives). We provide a proof for the sake of completeness.
@@
Let $\mathcal{G}=(\mathcal{A},(\Omega_{i})_{i\in\llbracket{}n\rrbracket{}})$ be a game such that $\Omega_{i}\in\{\mathsf{Reach}(T_{i}),\mathsf{B\ddot{u}chi}(T_{i}),\mathsf{coB\ddot{u}chi}(T_{i})\}$ for all $i\in\llbracket{}n\rrbracket{}$. Let $\pi=v_{0}v_{1}\ldots$ be a play. Then $\pi$ is the outcome of an NE from $v_{0}$ if and only if, for all $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$, $v_{\ell}\notin W_{i}(\Omega_{i})$ for all $\ell\in\mathbb{N}$.
@@
We let $\sigma=(\sigma_{i})_{i\in\llbracket{}n\rrbracket{}}$. By construction, we obtain that $\pi=\mathsf{Out}(\sigma,v_{0})$. To conclude, we show that $\sigma$ is an NE from $v_{0}$. Players in $\mathsf{SatPl}_{\mathcal{G}}(\pi)$ cannot have profitable deviations. It therefore remains to show that the other players do not have profitable deviations. We let $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$. It suffices to show that for all $\pi^{\prime}\in\mathsf{Plays}(\mathcal{A})$ starting in $v_{0}$, if $\pi^{\prime}$ is consistent with $\sigma_{-i}$, then $\pi^{\prime}\notin\Omega_{i}$. Let $\pi^{\prime}$ be a play starting in $v_{0}$ that is consistent with $\sigma_{-i}$. It follows from the definition of $\sigma$ that $\pi^{\prime}$ can be written as the combination of a prefix $\pi{\ell}$ of $\pi$ and a play $\pi^{\prime\prime}$ consistent with $\tau_{-i}$ starting in the winning region $W_{-i}(\mathsf{Plays}(\mathcal{A})\setminus\Omega_{i})$ of $\mathcal{P}_{-i}$ in $\mathcal{G}_{i}$. In particular, we have $\pi^{\prime\prime}\notin\Omega_{i}$. If $\Omega_{i}$ is a reachability objective, the target of $\mathcal{P}_{i}$ does not occur in $\pi{\ell}$ (as $\pi\notin\Omega_{i}$) and does not occur in $\pi^{\prime\prime}$ either (as $\pi^{\prime\prime}\notin\Omega_{i}$). If $\Omega_{i}$ is a Büchi or co-Büchi objective, we have $\pi^{\prime}\in\Omega_{i}$ if and only if $\pi^{\prime\prime}\in\Omega_{i}$. In both cases, we obtain that $\pi^{\prime}\notin\Omega_{i}$. This shows that $\mathcal{P}_{i}$ does not have a profitable deviation, and thus that $\sigma$ is an NE from $v_{0}$.
For safety objectives, we have to slightly adapt the characterisation of Theorem 4.1. If a safety objective is violated by a play, this violation is witnessed by a prefix of the play. It follows that there can be no profitable deviations for some $\mathcal{P}_{i}$ that occur at a point of an outcome after a vertex unsafe for $\mathcal{P}_{i}$ is traversed. In other words, in a safety game $\mathcal{G}=(\mathcal{A},(\mathsf{Safe}(T_{i}))_{i})$, a play $\pi$ is an NE outcome if and only if, for all $i\in\llbracket{}n\rrbracket{}\setminus\mathsf{SatPl}_{\mathcal{G}}(\pi)$, all vertices of $\pi$ that occur before $T_{i}$ is visited are outside of $W_{i}(\mathsf{Safe}(T_{i}))$. We omit a proof of this result: it can be shown in the same way as the corresponding finite-arena characterisation of [16] (or with an argument analogous to that of Theorem 4.1).